Area bounded by x^2=y, x=1 and x-axis.


thanks

Answer by Alan3354(21581)

Area bounded by x^2=y, x=1 and x-axis.
------------------
y = x^2 passes thru the Origin, so it's the area under the curve from x = 0 to x = 1.
----------
Integral(x^2) = x^3/3 + C
Area = f(1) - f(0)
= 1/3 - 0
Area = 1/3 square units